OpenStudy (anonymous):

how do i solve this when there is a ^(1/2)? Please help

OpenStudy (anonymous):

at x^1/2 I am guessing it is downward sloping.

OpenStudy (anonymous):

te slope can actually be calculated using calcuus

OpenStudy (anonymous):

dU/dx = nx ^n-1 1/2 (x) 1/2/-1 1/2 (x)^ -1/2

OpenStudy (anonymous):

@jim_thompson5910 can you check my figures jim?

jimthompson5910 (jim_thompson5910):

That looks good, if U(x) = x^(1/2), then U ' (x) = (1/2)x^(-1/2)

OpenStudy (anonymous):

slope m = 1/2a ^(-1/2)